# Detecting stochastic gravitational wave background from cosmic strings with next-generation detector networks: Component separation based on a multi-source astrophysical foreground noise model

### v1

@@ Abstract
Detecting stochastic gravitational wave background (SGWB) from cosmic strings is crucial for unveiling the evolutionary laws of the early universe and validating non-standard cosmological models. This study presents the first systematic evaluation of the detection capabilities of next-generation ground-based gravitational wave detector networks for cosmic strings. By constructing a hybrid signal model incorporating multi-source astrophysical foreground noise, including compact binary coalescences (CBCs) and compact binary hyperbolic encounters (CBHEs), we propose an innovative parameter estimation methodology based on multi-component signal separation. Numerical simulations using one-year observational data reveal three key findings: (1) The CE4020ET network, comprising the Einstein Telescope (ET-10 km) and the Cosmic Explorer (CE-40 km and CE-20 km), achieves nearly one order of magnitude improvement in constraining the cosmic string tension $G\mu$ compared to individual detectors, reaching a relative uncertainty $\Delta G\mu/G\mu<0.5$ for $G\mu>3.5\times 10^{-15}$ under standard cosmological framework; (2) The network demonstrates enhanced parameter resolution in non-standard cosmological scenarios, providing a novel approach to probe pre-Big Bang Nucleosynthesis cosmic evolution; (3) Enhanced detector sensitivity amplifies CBHE foreground interference in parameter estimation, while precise modeling of such signals could further refine $G\mu$ constraints by $1-2$ orders of magnitude. This research not only quantifies the detection potential of third-generation detector networks for cosmic string models but also elucidates the intrinsic connection between foreground modeling precision and cosmological parameter estimation accuracy, offering theoretical foundations for optimizing scientific objectives of next-generation gravitational wave observatories.
@@ I Introduction
SGWBs originate from various sources, which can be broadly classified into cosmological and astrophysical sources. Cosmological sources include cosmic strings (CSs) Nielsen and Olesen (1973); Kibble (1976); Vilenkin (1985); Witten (1985); Hindmarsh and Kibble (1995); Polchinski (2004), first-order phase transitions Krauss (1992); Kosowsky et al. (1992a, b); Kosowsky and Turner (1993); Kamionkowski et al. (1994); Hindmarsh et al. (2014), inflation Grishchuk (1975); Starobinsky (1979); Starobinskii (1979); Rubakov et al. (1982); Grishchuk (1993); Barnaby et al. (2012), and other early universe phenomena. Astrophysical sources include compact-binary coalescences (CBCs) (including BBHs, NSBHs, BNSs and white dwarf binaries) Meacher et al. (2015); Farmer and Phinney (2003), compact-binary hyperbolic encounters (CBHEs) Capozziello et al. (2008); De Vittori et al. (2012); García-Bellido and Nesseris (2018); Garcia-Bellido and Nesseris (2017); García-Bellido et al. (2022), stellar core collapse Ferrari et al. (1999); Buonanno et al. (2005); Crocker et al. (2015, 2017); Finkel et al. (2022) and others. These SGWB signals carry crucial information about cosmic structure formation and evolutionary processes, offering unprecedented insights into diverse physical phenomena, from compact object dynamics to early universe evolution.
@@
Detecting the SGWB from CSs is challenging because the GW signals from CSs are weak and obscured by intense astrophysical foreground noise dominated by CBCs. Second-generation detectors, including advanced LIGO Harry (2010); Aasi et al. (2015), advanced Virgo Acernese et al. (2015), and KAGRA Aso et al. (2013), have not yet detected any cosmological SGWB signals. This is due to their limited sensitivity and the inability to resolve most of the CBC foreground Abbott et al. (2023, 2021a). With the deployment of next-generation (NG) observatories, we expect to observe a more comprehensive SGWB of cosmological origin. The NG observatory network is expected to detect hundreds of thousands of CBC events annually Evans et al. (2021); Borhanian and Sathyaprakash (2024); Ronchini et al. (2022); Iacovelli et al. (2022); Evans et al. (2023), as well as many CBHE events Mukherjee et al. (2021). These observations will enable the development of more precise astrophysical foreground models, facilitating the detection of the SGWB from CSs.
In this study, we explore the ability of NG ground-based GW detector networks to detect CSs. Our objective is to determine whether ground-based GW detectors can detect non-standard features in the SGWB spectrum from CSs, thereby contributing to the reconstruction of the universe’s pre-BBN history. We employ the component separation method Parida et al. (2016) to isolate CS signals from astrophysical foregrounds. Our analysis includes both the standard cosmological model for CSs and the impact of non-standard cosmologies on their SGWB spectrum. We consider a more comprehensive range of astrophysical foreground noise, including both CBC and CBHE foregrounds. Additionally, we investigate the impact of excluding the CBHE foreground on the precision of CS constraints.
@@ II Sensitivity for detector networks
Figure 1: Sensitivity of different detector networks in detecting the SGWB. Left: Amplitude spectral densities for the three detectors in the networks. Middle: Overlap reduction functions for each baseline in different detector networks, normalized such that $\gamma(f)=1$ for co-located and co-aligned L-shaped detectors. Right: $1\sigma$ PI curves showing the sensitivity of the various detector networks considered in this study to the stochastic gravitational-wave background Thrane and Romano (2013).
@@ III.1 Stochastic gravitational wave background from cosmic strings
CSs are one-dimensional topological defects that arise from the spontaneous breaking of U(1) symmetry in the early universe Kibble (1976). The primary source of GWs emitted by a CS network is the continuous formation of loops during network fragmentation, occurring when strings self-intersect or collide. These loops, acting as sources of GWs, persist throughout the entire history of the universe, contributing to a SGWB through the superposition of numerous uncorrelated emissions Vilenkin (1981); Vachaspati and Vilenkin (1985); Damour and Vilenkin (2000, 2005); Siemens et al. (2007); Olmez et al. (2010). The impact of cosmology on the GW spectrum arises from two competing effects: the red-tilt (caused by the redshifting of GW energy density) and the blue-tilt (due to the loop production rate). During radiation domination, these two effects cancel each other out, leading to a flat spectrum Gouttenoire et al. (2020). Since the frequency range of ground-based GW detectors corresponds to the CS GW spectrum during the radiation era, studying the characteristics of this spectrum can provide valuable insights into the universe’s expansion history.
@@ III.1.1 Standard cosmological model
×$\times\left\{\frac{\left({\Omega_{m}}/{\Omega_{r}}\right)^{1/4}}{\left(B_{m}\left({\Omega_{m}}/{\Omega_{r}}\right)^{1/2}+f\right)^{1/2}}\left[2+\frac{f}{B_{m}\left({\Omega_{m}}/{\Omega_{r}}\right)^{1/2}+f}\right]-\frac{1}{(B_{m}+f)^{1/2}}\left[2+\frac{f}{B_{m}+f}\right]\right\},$ (4)
@@ III.1.2 Non-standard cosmological model
where $\text{where }d=$$\begin{cases}1&\text{for }w\leq\frac{1}{9},\\ -2\frac{3w-1}{3w+1}&\text{for }w>\frac{1}{9}.\end{cases}$ (8)
@@
in this case, we do not consider the effect of the effective relativistic DoF and set $T_{\rm rd}=1\rm GeV$. For values of $w$ ranging from $w=1/9$ to $w=1$ , the spectrum evolves differently. For $w\leq 1/9$, the predicted spectrum remains unchanged, closely matching the spectrum of strings formed during inflation Guedes et al. (2018). For $31/9<w<1/$ 3, the spectrum exhibits a negative slope for $f\geq f_{\rm rd}$. When $w=1/3$, the spectrum corresponds to that of the standard cosmological model. For $w>1/3$, the spectrum exhibits a positive slope for $f\geq f_{\rm rd}$. When $w=1$, the results align with the spectrum of CSs during the kination-dominated era immediately following inflation Gouttenoire et al. (2020). Additionally, Ref. Zhu et al. (2024) analyzes a combination of pulsar timing array (PTA), BBN, and cosmic microwave background data, yielding $w=0.44^{+0.52}_{-0.40}$ at the 95% confidence level. Therefore, we set $w=1$, $w=1/2$, $w=1/3$, and $w=1/4$ to investigate these different scenarios.
@@
where $f_{\ast}=25\rm Hz$, $\Omega_{\rm CS}=\Omega_{\rm CS}(f_{\ast})$ as defined in Eq. (6), and $\alpha_{\rm CS}=-d$ as defined in Eq. (8).
@@ III.3 Stochastic gravitational wave background from compact-binary hyperbolic encounters
$\Omega_{\rm CBHE}(f)=\frac{n_{\rm gc}}{V\rho_{\rm c}}\int^{\infty}_{0}\frac{\mathcal{N}(z)\mathcal{K}(z)dz}{(1+z)^{3}H(z)},$ (13)
where $n_{\rm gc}$ is the number of GCs per Milky Way-equivalent galaxy (MWEG), $V$ is the comoving volume up to redshift $z$, $\mathcal{N}(z)$ is the number of MWEGs between redshift $z$ and $z+{\rm d}z$ as given in Eq. (32) of Ref. Kerachian et al. (2024), $\mathcal{K}(z)$ represents the probability per unit time of encounters occurring within each GC as given in Eq. (36) of Ref. Kerachian et al. (2024), $\rho_{\rm c}$ is the energy density of the universe, and $H(z)$ is the Hubble parameter at redshift $z$.
@@ IV.1 Component separation
×$\times\sum_{I,t,f}\gamma_{I}(f)\frac{\mathcal{F}^{\alpha}(f)\widetilde{s}_{I,1,t}^{\ast}(f)\widetilde{s}_{I,2,t}(f)}{f^{3}P_{I,1,t}(f)P_{I,2,t}(f)},$ (32)
@@ IV.2 Simulation data
We consider three data cases for the GW fractional energy density spectrum $\Omega_{\rm GW}$ in each detector:
1. $\Omega_{\rm GW}(f)=\Omega_{\rm CBC}(f)+\Omega_{\rm CBHE}(f)$, that is, two components: only astrophysical foreground noise.
2. $\Omega_{\rm GW}(f)=\Omega_{\rm CS}(f)+\Omega_{\rm CBC}(f)+\Omega_{\rm CBHE}(f)$, that is, three components: CS GW signals and astrophysical foreground noise. In this case, the CBHE component is omitted during signal separation, and we investigate how proper modeling of the CBHE component improves CS detection.
3. $\Omega_{\rm GW}(f)=\Omega_{\rm CS}(f)+\Omega_{\rm CBC}(f)+\Omega_{\rm CBHE}(f)$, that is, three components: like data case 2, but consider both the CBC and CBHE foregrounds in full.
@@ V Results
$\log\mathcal{L}(\hat{\Omega}|\theta)=$ $-\frac{1}{2}\left[\left(\Omega(\theta)-\hat{\Omega}\right)^{\rm T}\Sigma^{-1}\left(\Omega(\theta)-\hat{\Omega}\right)\right.$
@@
Data Case 2 $1$ $6.3\times 10^{-16}$ $2.0\times 10^{-15}$ $5.6\times 10^{-15}$
@@
$1/4$ $>1.0\times 10^{-10}$ $>1.0\times 10^{-10}$ $>1.0\times 10^{-10}$
Data Case 3 $1$ $2.2\times 10^{-16}$ $6.3\times 10^{-17}$ $2.8\times 10^{-17}$
@@
$1/3$ $1.4\times 10^{-13}$ $1.1\times 10^{-14}$ $3.5\times 10^{-15}$

### v2

@@ Abstract
Detecting stochastic gravitational wave background (SGWB) from cosmic strings is crucial for unveiling the evolutionary laws of the early universe and validating non-standard cosmological models. In light of extensive existing studies on the detection capabilities of third-generation gravitational wave detectors for cosmic string signals, this work provides a systematic evaluation focused specifically on the performance of next-generation ground-based detector networks. By constructing a hybrid signal model that incorporates multi-source astrophysical foregrounds, including compact binary coalescences (CBCs) and compact binary hyperbolic encounters (CBHEs), we develop a parameter estimation methodology based on multi-component signal separation. Numerical simulations using one-year observational data reveal three key findings: (1) The CE4020ET network, comprising the Einstein Telescope (ET-10 km) and the Cosmic Explorer (CE-40 km and CE-20 km), achieves nearly one order of magnitude improvement in constraining the cosmic string tension $G\mu$ compared to individual detectors, reaching a relative uncertainty $\Delta G\mu/G\mu<0.5$ for $G\mu>3.5\times 10^{-15}$ under standard cosmological framework; (2) The network demonstrates enhanced parameter resolution in non-standard cosmological scenarios, providing a novel approach to probe pre-Big Bang Nucleosynthesis cosmic evolution; (3) Enhanced detector sensitivity amplifies CBHE foreground interference in parameter estimation, while precise modeling of such signals could further refine $G\mu$ constraints by 121-2 orders of magnitude. This research not only quantifies the detection potential of third-generation detector networks for cosmic string models but also elucidates the intrinsic connection between foreground modeling precision and cosmological parameter estimation accuracy, offering theoretical foundations for optimizing scientific objectives of next-generation gravitational wave observatories.
@@ I Introduction
SGWBs originate from various sources, which can be broadly classified into cosmological and astrophysical sources. Cosmological sources include cosmic strings (CSs) Nielsen and Olesen (1973); Kibble (1976); Vilenkin (1985); Witten (1985); Hindmarsh and Kibble (1995); Polchinski (2004), first-order phase transitions Krauss (1992); Kosowsky et al. (1992a, b); Kosowsky and Turner (1993); Kamionkowski et al. (1994); Hindmarsh et al. (2014), inflation Grishchuk (1975); Starobinsky (1979); Starobinskii (1979); Rubakov et al. (1982); Grishchuk (1993); Barnaby et al. (2012), and other early universe phenomena. Astrophysical sources include compact-binary coalescences (CBCs) (including BBHs, NSBHs, BNSs and white dwarf binaries) Meacher et al. (2015); Farmer and Phinney (2003), compact-binary hyperbolic encounters (CBHEs) Capozziello et al. (2008); De Vittori et al. (2012); García-Bellido and Nesseris (2018); Garcia-Bellido and Nesseris (2017); García-Bellido et al. (2022), stellar core collapse Ferrari et al. (1999); Buonanno et al. (2005); Crocker et al. (2015, 2017); Finkel et al. (2022) and others. These SGWB signals are expected to carry crucial information about cosmic structure formation and evolutionary processes, offering unprecedented insights into diverse physical phenomena ranging from compact object dynamics to early universe evolution.
@@
Detecting the SGWB from cosmic strings poses significant challenges, as the corresponding gravitational wave signals are inherently weak and expected to be obscured by intense astrophysical foreground noise dominated by compact binary coalescences (CBCs). Second-generation detectors, including Advanced LIGO Harry (2010); Aasi et al. (2015), Advanced Virgo Acernese et al. (2015), and KAGRA Aso et al. (2013), have not yet confirmed the detection of any cosmological SGWB signals, with current results providing only upper limits. This is primarily attributed to their limited sensitivity and the present lack of effective methodologies to sufficiently resolve and subtract the predicted CBC foreground contamination Abbott et al. (2023, 2021a). With the deployment of next-generation (NG) observatories, we expect to observe a more comprehensive SGWB of cosmological origin. The NG observatory network is expected to detect hundreds of thousands of CBC events annually Evans et al. (2021); Borhanian and Sathyaprakash (2024); Ronchini et al. (2022); Iacovelli et al. (2022); Evans et al. (2023), as well as many CBHE events Mukherjee et al. (2021). These observations will enable the development of more precise astrophysical foreground models, facilitating the detection of the SGWB from CSs.
Building upon existing research on cosmic string detection with third-generation gravitational wave detectors Martinovic et al. (2021); Branchesi et al. (2023); Meijer et al. (2024), this study further investigates the joint detection capabilities of NG ground-based gravitational wave detector networks for CSs. Our objective is to systematically evaluate the networks’ ability to detect non-standard features in the SGWB spectrum generated by cosmic strings under conditions involving more complex astrophysical foreground noise, thereby providing new avenues for reconstructing the evolutionary history of the universe prior to BBN. We employ the component separation method Parida et al. (2016) to isolate CS signals from astrophysical foregrounds. Our analysis includes both the standard cosmological model for CSs and the impact of non-standard cosmologies on their SGWB spectrum. We consider a more comprehensive range of astrophysical foreground noise, including both CBC and CBHE foregrounds. Additionally, we investigate the impact of excluding the CBHE foreground on the precision of CS constraints.
@@ II Sensitivity for detector networks
Figure 1: Sensitivity of different detector networks in detecting the SGWB. Left: Amplitude spectral densities for the three detectors in the networks. Middle: Overlap reduction functions for each baseline in different detector networks, normalized such that $\gamma(f)=1$ for co-located and co-aligned L-shaped detectors. Right: $1\sigma$ PI curves showing the sensitivity of the various detector networks considered in this study to the stochastic gravitational-wave background after one year of observation Thrane and Romano (2013).
@@ III.1 Stochastic gravitational wave background from cosmic strings
This work investigates the gravitational wave background generated by networks of local cosmic strings. Local cosmic strings arise from the spontaneous breaking of a local gauge symmetry and represent cosmologically significant topological defects predicted by well-motivated particle physics models Kibble (1976), such as supersymmetric grand unified theories (SUSY GUTs). It is important to note that cosmic strings may also originate from more complex symmetry-breaking scenarios Jeannerot et al. (2003) or manifest as fundamental strings in superstring theory. In contrast, global cosmic strings, formed through the breaking of global symmetries, exhibit distinctly different phenomenological features, including their gravitational coupling and radiation efficiency, compared to local cosmic strings Baeza-Ballesteros et al. (2024); Klaer and Moore (2017); Chang and Cui (2020). The primary source of GWs from a cosmic string network is the continuous formation of loops through network fragmentation, which occurs when strings self-intersect or collide. These loops oscillate relativistically and act as persistent sources of GWs throughout cosmic history, collectively generating a SGWB through the superposition of numerous uncorrelated emissions Vilenkin (1981); Vachaspati and Vilenkin (1985); Damour and Vilenkin (2000, 2005); Siemens et al. (2007); Olmez et al. (2010). The impact of cosmology on the GW spectrum stems from two competing effects: a red-tilt due to the redshifting of GW energy density, and a blue-tilt from the loop production rate. During radiation domination, these effects cancel, resulting in a nearly scale-invariant (flat) spectrum Gouttenoire et al. (2020). Since the frequency band of ground-based GW detectors probes the part of the CS-generated SGWB originating from the radiation-dominated era, characterizing this spectrum offers a powerful probe into the expansion history of the universe.
@@ III.1.1 Standard cosmological model
$\times\left\{\frac{\left({\Omega_{m}}/{\Omega_{r}}\right)^{1/4}}{\left(B_{m}\left({\Omega_{m}}/{\Omega_{r}}\right)^{1/2}+f\right)^{1/2}}\left[2+\frac{f}{B_{m}\left({\Omega_{m}}/{\Omega_{r}}\right)^{1/2}+f}\right]-\frac{1}{(B_{m}+f)^{1/2}}\left[2+\frac{f}{B_{m}+f}\right]\right\},$ (4)
@@ III.1.2 Non-standard cosmological model
$\text{where }d=$ $\begin{cases}1&\text{for }w\leq\frac{1}{9},\\ -2\frac{3w-1}{3w+1}&\text{for }w>\frac{1}{9}.\end{cases}$ (8)
@@
in this case, we do not consider the effect of the effective relativistic DoF and set $T_{\rm rd}=1\rm GeV$. For values of $w$ ranging from $w=1/9$ to $w=1$ , the spectrum evolves differently. For $w\leq 1/9$, the predicted spectrum remains unchanged, closely matching the spectrum of strings formed during inflation Guedes et al. (2018). For $1/9<w<1/3$, the spectrum exhibits a negative slope for $f\geq f_{\rm rd}$. When $w=1/3$, the spectrum corresponds to that of the standard cosmological model. For $w>1/3$, the spectrum exhibits a positive slope for $f\geq f_{\rm rd}$. When $w=1$, the results align with the spectrum of CSs during the kination-dominated era immediately following inflation Gouttenoire et al. (2020). Additionally, Ref. Zhu et al. (2024) analyzes a combination of pulsar timing array (PTA), BBN, and cosmic microwave background data, yielding $w=0.44^{+0.52}_{-0.40}$ at the 95% confidence level. Therefore, we set $w=1$, $w=1/2$, $w=1/3$, and $w=1/4$ to investigate these different scenarios.
@@
where $f_{\ast}=25\rm Hz$, $\Omega_{\rm CS}=\Omega_{\rm CS}(f_{\ast})$ as defined in Eq. (6), and $\alpha_{\rm CS}=-d$ as defined in Eq. (8). We note that our calculations in Eqs. (3)–(5), (8) and (9) rely on a simplified approach: we assume cosmic string loops emit gravitational waves mainly at their fundamental frequency. While the complete theory includes all harmonic modes (following an $n^{-4/3}$ spectrum), we only consider the fundamental mode in this work. This simplification is reasonable because including higher harmonics has been shown to have little effect on the high-frequency part of the gravitational wave background Blanco-Pillado and Olum (2017); Sousa et al. (2020). Our method is consistent with previous studies that use similar simplifications to make the analysis tractable Blanco-Pillado et al. (2024).
@@ III.3 Stochastic gravitational wave background from compact-binary hyperbolic encounters
$\Omega_{\rm CBHE}(f)=\frac{n_{\rm gc}}{V\rho_{\rm c}}\int^{\infty}_{0}\frac{G(z)\mathcal{K}(z)dz}{(1+z)^{3}H(z)},$ (13)
where $n_{\rm gc}$ is the number of GCs per Milky Way-equivalent galaxy (MWEG), $V$ is the comoving volume up to redshift $z$, $G(z)$ is the number of MWEGs between redshift $z$ and $z+{\rm d}z$ as given in Eq. (32) of Ref. Kerachian et al. (2024), $\mathcal{K}(z)$ represents the probability per unit time of encounters occurring within each GC as given in Eq. (36) of Ref. Kerachian et al. (2024), $\rho_{\rm c}$ is the energy density of the universe, and $H(z)$ is the Hubble parameter at redshift $z$.
@@ IV.1 Component separation
$\times\sum_{I,t,f}\gamma_{I}(f)\frac{\mathcal{F}^{\alpha}(f)\widetilde{s}_{I,1,t}^{\ast}(f)\widetilde{s}_{I,2,t}(f)}{f^{3}P_{I,1,t}(f)P_{I,2,t}(f)},$ (32)
@@ IV.2 Simulation data
To systematically investigate the impact of precise astrophysical foreground modeling on the detection capability of cosmic string signals, we designed three data scenarios in each detector. The key distinction among them lies in whether the signal components injected into the simulated data match the template model components used in Bayesian parameter estimation. The specific configurations are summarized below:
1. The injected signal model is given by $\Omega_{\rm GW}(f)=\Omega_{\rm CBC}(f)+\Omega_{\rm CBHE}(f)$, and the analysis template is chosen to be identical to the injection model. This configuration serves as a null test to validate the robustness of our methodology, specifically examining whether it can correctly infer the absence of cosmic string signals when only astrophysical foregrounds are present.
2. The injected signal model is given by $\Omega_{\rm GW}(f)=\Omega_{\rm CS}(f)+\Omega_{\rm CBC}(f)+\Omega_{\rm CBHE}(f)$, and the analysis template intentionally excludes the CBHE component. This configuration investigates the impact of incomplete foreground modeling by deliberately omitting the CBHE component during signal separation, allowing quantitative assessment of how proper inclusion of this foreground improves cosmic string detection sensitivity and parameter estimation accuracy.
3. The injected signal model is given by $\Omega_{\rm GW}(f)=\Omega_{\rm CS}(f)+\Omega_{\rm CBC}(f)+\Omega_{\rm CBHE}(f)$, and the analysis template is chosen to be identical to the injection model. This configuration demonstrates the optimal constraining power on cosmic string parameters under a complete and precise modeling framework.
@@ V Results
[…]
@@
Data Case 2 11 $6.3\times 10^{-16}$ $2.0\times 10^{-15}$ $5.6\times 10^{-15}$
@@
$1/4$ $\times 10^{-10} >1.0×10$$\times 10^{-10}$ $1.0\times 10^{-10}$
Data Case 3 11 $2.2\times 10^{-16}$ $6.3\times 10^{-17}$ $2.8\times 10^{-17}$
@@
$1/3$ $4\times 10^{-13} 1.1$$1\times 10^{-14}$ $3.5\times 10^{-15}$